. A woman exerts a force of 500 N straight up to lift a 35 kg basket. What is the acceleration of the basket?


1
Expert's answer
2021-09-17T15:33:02-0400

According to the second Newton's law, the acceleration is given as follows:


"a = \dfrac{F-mg}{m}"

where "F = 500N" is the exerted force, "mg" is the gravity force ('-' sign since exerted force and gravity force have opposite directions), "m = 35kg" is the mass of the basket, and "g = 9.8m\/s^2" is the gravitational acceleration. Thus, obtain:


"a = \dfrac{F}{m} - g = \dfrac{500N}{35kg} - 9.8\dfrac{m}{s^2} \approx 4.5 m/s^2"

Answer. 4.5 m/s^2.